A Bibliography of Publications of Roman Pozarlik

Roman Pozarlik Wroclaw University of Technology Institute of Engineering Cybernetics

Janiszewskiego Str. 11-17

PL 50-370 Wroclaw Poland

Tel: +48 71 320 20 15 FAX: +48 71 21 26 77

E-mail: rpoz@pwr.wroc.pl (Internet)

25 November 2011 Version 1.03

Abstract

This bibliography records publications of Roman Pozarlik.

Title word cross-reference

/ [Ano94].

4th [BGH98].

8th [NBZ98].

Alaska [IEE98]. Anchorage [IEE98]. Annual [Ano94]. April [BGH98]. Artificial [NBZ98]. August [SL97].

CA [SL97]. California [Ano94]. Causal [Poz97]. Cognitive [Poz97]. Computation [Poz95]. Computational [IEE98].

Computations [Poz98b]. conference [SL97]. Congress [Ano94]. Connectionist [Poz97]. Country [Ano94].

DC [Ano95]. Diego [Ano94]. Distributed [Poz94]. Do [Poz98c].

Finite [Poz98c].

Hotel [Ano94].

ICANN [NBZ98]. IEEE [IEE98]. Integrating [Poz94]. Intelligence [IEE98]. International [Ano94].

Jigsaw [Poz94]. Joint [IEE98]. July [Ano95]. June [Ano94].

London [BGH98].

May [IEE98]. Meeting [Ano94].

Network [Poz94]. Networks [Poz98b]. Neural [Poz94]. nineteenth [SL97].

Parallel [Poz94]. Perform [Poz98c]. Proceedings [SL97]. Processing [Poz94]. Psychology [BGH98].

Recurrent [Poz98c]. Renaissance [Ano95]. Representation [Poz95]. Representations [Poz97].

San [Ano94]. Science [SL97]. September [NBZ98]. Sequential [Poz98b]. Skövde [NBZ98]. Society [Ano94]. Stanford [SL97]. Subsymbolic [Poz95]. Sweden [NBZ98]. Symbol [Poz98a]. Symbolic [Poz94].

Town [Ano94]. Type [Poz98c].

University [SL97]. Unstructured [Poz95]. USA [Ano94].

Washington [Ano95]. WCNN’95 [Ano95]. Workshop [BGH98]. World [Ano94].

References

Anonymous:1994:WCN

[Ano94] Anonymous, editor. World Congress on Neural Networks, San Diego / 1994 International Neural Network Society Annual Meeting, Town and Country Hotel, San Diego, California, USA, June 5–9, 1994. Lawrence Erlbaum Associates, Hillsdale, NJ, 1994. ISBN 0-8058-1745-X. LCCN QA76.87.W68 1994. Four volumes.

Anonymous:1995:WWD

[Ano95] Anonymous, editor. WCNN’95, Washington, DC: World Congress on Neural Networks: 1995 International Neural Network Society Annual Meeting: Renaissance Hotel, Washington, DC, USA, July 17–21, 1995. Lawrence Erlbaum Associates, Hillsdale, NJ, 1995. ISBN 0-8058-2125-2. LCCN QA76.87.W68 1995. Two volumes.

Bullinaria:1998:NCP

[BGH98] John A. Bullinaria, David W. Glasspool, and George Houghton, editors. 4th Neural Computation and Psychology Workshop: London, 9–11 April 1997: connectionist representations, Perspectives in neural computing. Springer-Verlag, London, UK, 1998. ISBN 3-540-76208-6. LCCN QA76.87.N473 1997.

IEEE:1998:IIJ

[IEE98] IEEE, editor. The 1998 IEEE International Joint Conference on Neural Networks: proceedings: IEEE World Congress on Computational Intelligence: May 4–May 9, 1998, Anchorage, Alaska, USA. IEEE Press, Piscataway, NJ, 1998. ISBN 0-7803-4859-1 (softbound), 0-7803-4860-5 (casebound), 0-7803-4861-3 (microfiche), 0-7803-4862-1 (CD-ROM). LCCN QA76.87.I57 1998. Three volumes. Also known as WCCI 98 and IJCNN ’98. IEEE Catalog Number 98CH36227.

Niklasson:1998:IPI

[NBZ98] Lars F. Niklasson, Mikael B. Boden, and T. (Tom) Ziemke, editors. ICANN 98: proceedings of the 8th International Conference on Artificial Neural Networks, Skövde, Sweden, 2–4 September 1998. Springer-Verlag, London, UK, 1998. ISBN 3-540-76263-9. LCCN QA76.87.I56 1998. Two volumes.

Pozarlik:1994:ISP

[Poz94] Roman Pozarlik. Integrating symbolic and parallel distributed processing in a jigsaw neural network. In Anonymous [Ano94], pages 550– 555. ISBN 0-8058-1745-X. LCCN QA76.87.W68 1994. Four volumes.

Pozarlik:1995:URS

[Poz95] Roman Pozarlik. An unstructured representation for subsymbolic computation. In Anonymous [Ano95], pages 309–312. ISBN 0-8058-2125-2. LCCN QA76.87.W68 1995. Two volumes.

Pozarlik:1997:CCC

[Poz97] Roman Pozarlik. Cognitive computation on connectionist causal representations. In Shafto and Langley [SL97], pages 1013–?? ISBN 0-8058-2941-5. LCCN ????

Pozarlik:1998:CSP

[Poz98a] Roman Pozarlik. Connectionist symbol processing with causal representations. In Bullinaria et al. [BGH98], pages 331–342. ISBN 3-540-76208-6. LCCN QA76.87.N473 1997.

Pozarlik:1998:IPC

[Poz98b] Roman Pozarlik. Integrating parallel computations and sequential symbol processing in neural networks. In IEEE [IEE98], pages 1444–1449. ISBN 0-7803-4859-1 (softbound), 0-7803-4860-5 (casebound), 0-7803-4861-3 (microfiche), 0-7803-4862-1 (CD-ROM). LCCN QA76.87.I57 1998. Published in CD version.

Pozarlik:1998:WTF

[Poz98c] Roman Pozarlik. What type of finite computations do recurrent neural networks perform? In Niklasson et al. [NBZ98], pages 1065– 1070. ISBN 3-540-76263-9. LCCN QA76.87.I56 1998. Two volumes.

Shafto:1997:PNA

[SL97] Michael G. Shafto and Pat. Langley, editors. Proceedings of the nineteenth annual conference of the Cognitive Science Society: August 7–10, 1997, Stanford University, Stanford, CA. Lawrence Erlbaum Associates, Hillsdale, NJ, 1997. ISBN 0-8058-2941-5. LCCN ????